Saul Goodman, Jesse Pinkman, and Walter White formed a partnership on January 1, 2020, and made capital contributions of $125,000 (Goodman), $175,000 (Pinkman), and $250,000 (White), respectively. With respect to the division of income, they agreed to the following:

 (1) interest of an amount equal to 10% of the that partner’s beginning capital balance for the year; 

(2) annual compensation of $15,000 to Pinkman; and (3) the remainder of the income or loss to be split among the partners in the following percentages: 

(a) 20% for Goodman; 

(b) 40% for Pinkman; and 

(c) 40% for White. 

Net income was $200,000 in 2020 and $240,000 in 2021. Each partner withdrew $1,500 for personal use every month during 2020 and 2021. What was White’s total share of net income for 2021?

            A)    $99,060. 
            B)    $126,900.
            C)    $62,160.
            D)    $93,060.
            E)    $96,000.
Answer #1

> Brilliant work mate! Thank you for your help.

Jack Who Fri, Aug 13, 2021 7:48 AM
Answer #2

What was White’s capital balance at the end of 2020?


White's capital
Beginning capital250000
Add: Share in net income77000
Less: Drawings-1500*12= -18000


Balance at end of 2020309000

Working :

Share in net income :


GoodmanPinkmanWhiteTotal
Interest on capital125000*10%=12500175000*10%=17500250000*10%=2500055000
Salary allowance
15000
15000
Remaining income [200000-55000-15000=130000]130000*20%=26000130000*40%=5200052000130000





Total share in net income for year 2020385008450077000200000


answered by: studymoon

> thank you, this was part one

Jack Who Fri, Aug 13, 2021 7:49 AM